How to find probability distribution for example like car sale and probability. So I need a method to find out number of sales in two day period.

1
Expert's answer
2022-02-07T17:23:43-0500

You can find probability distribution using the method shown in below example.


Example: Construct a probability distribution of the random variable X for the number of false answers in a 4-item true or false test.


Solution:

X = the number of false answers

X = {0,1,2,3,4}

"p=q=\frac 12,n=4"

"P(X=0)=^4C_0(\frac 12)^0(\frac 12)^4=\frac 1{16}
\\P(X=1)=^4C_1(\frac 12)^1(\frac 12)^3=\frac 4{16}=\frac 14
\\ P(X=2)=^4C_2(\frac 12)^2(\frac 12)^2=\frac 6{16}=\frac 38
\\ P(X=3)=^4C_3(\frac 12)^3(\frac 12)^1=\frac 4{16}=\frac 14
\\P(X=4)=^4C_4(\frac 12)^4(\frac 12)^0=\frac 1{16}"

Thus, probability distribution of X is:

X P(X)

0 1/16

1 1/4

2 3/8

3 1/4

4 1/16